OpenStudy (benito893764):

find the area of a regular pentagram if one of the sides of inner pentagon is 3cm

OpenStudy (wolf1728):

To find the area of a pentagon we use the formula (where s = side): https://en.wikipedia.org/wiki/Pentagon AREA = s^2 * square root (25 + 10 * sqrt(5)) /4 AREA = 9 * sq root (25 + 22.361) / 4 AREA = 9 * 6.8819 / 4 AREA = 61.9371 / 4 PENTAGON AREA = 15.484 The pentagon is surrounded by 5 isosceles triangles whose height is calculated by tan (72 degrees) = height / 1.5 height = 3.0777 * 1.5 height = 4.6166 and the area of EACH triangle equals 1.5 * 4.6166 = 6.9249 AREA of all 5 triangles = 34.625 PLUS PENTAGON AREA = 15.484 TOTAL PENTAGRAM AREA = 50.109

OpenStudy (benito893764):

thanks ^_^